OpenStudy (anonymous):

Choose the symbol that makes the following sentence true: 1 3/4 ____ 1.75%. A.< B.> C.=

OpenStudy (anonymous):

C.

Directrix (directrix):

1 3/4 = 1.75 but not 1.75%. Did you mean to write 1.75% in the question?

OpenStudy (anonymous):

3/4=.75 both whole numbers are the same so they are equal

Directrix (directrix):

1.75% = 0.0175

Directrix (directrix):

True: 1 3/4 = 1.75 False: 1.75 = 1.75% 1.75% means 1.75 over 100 because percent means "hundredths."

OpenStudy (anonymous):

ummm didnt see the percent

Directrix (directrix):

Answer: B.>

OpenStudy (radar):

don't like the options, but again it is not my problem. I think they should of left off % the percent symbol so that C would of been correct.